On a frictionless horizontal air table, puck A (wi
Answer #1

Part A :

According to the conservation of momentum,
m1v1 + m2v2 = m1v3 + m2v4
Set right as positive
0.247v1 = 0.247*(-0.120) + 0.369 *0.648
v1 = 0.848 ms^-1


Part B:

Initial kinetic energy = 0.247 *0.848^2 /2 = 0.178 J
Final kinetic energy = 0.247 *0.120^2 /2 + 0.369 *0.648^2 /2

= 0.0793 J
Thus, the total kinetic energy has decreased.